OpenStudy (anonymous):

A right triangle’s legs are 3 and 4 meters long. What is the measure of the angle adjacent to the 4-meter leg to the nearest tenth of a degree?

OpenStudy (anonymous):

\[\frac{180 \text{ArcTan}\left[\frac{3}{4}\right]}{\pi }=36.9{}^{\circ} \]

OpenStudy (anonymous):

\[\tan^{-1} (3/4)\]